OpenStudy (coolshubhs):

make an equation for this.. : Two years ago,a man's age was three times the square of his son's age.In three years time, his age will be four times his son's age.Find their present ages.

OpenStudy (anonymous):

in one or two unknowns?

OpenStudy (coolshubhs):

it's given find their present ages..

OpenStudy (anonymous):

sorry had no connection ok if x is the current age of the father and y that of the son then: two years ago the father was (x-2) and the son was (y-2) =>(x-2)=3(y-2)^2........equation 1 and in three years the father will be (x+3) and the son (y+3) =>(x+3)=4(y+3)....equation 2 now try to solve the system for x and y

OpenStudy (coolshubhs):

isn't is possible to make an equation with one variable?

OpenStudy (anonymous):

it is,but sounds a little weird let x be the son's age in 3 years his father's age will be 4x two years ago the son was x-5 and the father was 4x-5 =>(4x-5)=3(x-5)^2

OpenStudy (coolshubhs):

i made an equation like: 3(x-2)^2=4(x+3) but it's not giving the right answer.

OpenStudy (anonymous):

the answers should be 29 and 5

OpenStudy (coolshubhs):

wow great.. it's right.. which equation you used??

OpenStudy (anonymous):

the system in x and y is actually easier to understand if you use my second approach you'll get the age of the son in 3 years,so you have to subtract 3 from x...

OpenStudy (coolshubhs):

but i think i have to do it in single variable equation.. just tell me what's the correct equation for this?